OpenStudy (howard-wolowitz):

Given two terms from a geometric sequence, identify the first term and the common ratio:

OpenStudy (howard-wolowitz):

\[a _{5}=48 .... a _{11} = 3072\]

OpenStudy (howard-wolowitz):

Hey!

OpenStudy (howard-wolowitz):

A. \[a _{1} =2 ......... r=3\]

OpenStudy (howard-wolowitz):

B. \[a _{1}=2 ...... r=2\]

OpenStudy (howard-wolowitz):

C. \[a _{1}=3 ...... r=3\]

OpenStudy (howard-wolowitz):

D. \[a _{1}=3....... r=2\]

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (howard-wolowitz):

Hey I got 15 questions and I need your help if you will help me I'll medal you each time ok!

OpenStudy (michele_laino):

please us this formula: \[{a_{11}} = {a_1}{q^{10}}\] where q is the common ratio

OpenStudy (michele_laino):

and this one: \[{a_5} = {a_1}{q^4}\]

OpenStudy (michele_laino):

if I divide side bu side those above formulas, I get: \[\frac{{{a_{11}}}}{{{a_5}}} = {q^6}\] now substitute your data, and you will get the value for q

OpenStudy (michele_laino):

oops...side by side...

OpenStudy (michele_laino):

what do you get?

OpenStudy (howard-wolowitz):

Ok, I gotcha and I got B.

OpenStudy (michele_laino):

hint: \[\frac{{{a_{11}}}}{{{a_5}}} = \frac{{3072}}{{48}} = 64\]

OpenStudy (howard-wolowitz):

ok but thats not an answer

OpenStudy (howard-wolowitz):

am i supposed to combine that some how

OpenStudy (michele_laino):

please your common ratio q is: \[64 = {q^6}\] so \[\begin{gathered} q = 2 \hfill \\ and\;\qquad\\ {a_1} = \frac{{{a_5}}}{{{q^4}}} = \frac{{48}}{{16}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (howard-wolowitz):

ohhh so 3

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

because of 6 and 2

OpenStudy (howard-wolowitz):

ok i gotcha

OpenStudy (howard-wolowitz):

so r =3

OpenStudy (howard-wolowitz):

and a1 equals 3 so C.

OpenStudy (michele_laino):

I think q=r = 2 since 64= 2^6

OpenStudy (howard-wolowitz):

kk d next question

OpenStudy (michele_laino):

so your answer is option D.